OpenStudy (anonymous):

5=y+x into slope intercept form

OpenStudy (anonymous):

solve for y y = 5-x

Parth (parthkohli):

To get the slope-intercept form, just solve for \(\rm y\)

Parth (parthkohli):

@jayz657 No, not exactly that.

Parth (parthkohli):

You want the \(\rm x \) term first.

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

??? i dont see whats wrong lol your slope is -1 and intercept is 5 is that wrong?

OpenStudy (anonymous):

is it y=- 5-x

OpenStudy (anonymous):

does that look like y=mx+b your m=-1 and b=5 so now write it

OpenStudy (anonymous):

lol you want it like that? it still the same as my answer though but ok it is y=-x+5

OpenStudy (anonymous):

ok